OpenStudy (bloomlocke367):

How do you find the determinant of a 3 x 3 matrix?

OpenStudy (bloomlocke367):

@rational

OpenStudy (rational):

\[\begin{vmatrix} a&b&c\\d&e&f\\g&h&i \end{vmatrix}\]

OpenStudy (bloomlocke367):

um, what? XD

OpenStudy (rational):

first ignore first row and first column : \[\begin{vmatrix} \color{gray}{ a}&\color{gray}{ b}&\color{gray}{ c} \\\color{gray}{d}&e&f\\\color{gray}{g}&h&i \end{vmatrix} \] the first term of determinant would be \(a(ei-hf)\)

OpenStudy (rational):

did u watch that video

OpenStudy (bloomlocke367):

yes

OpenStudy (bloomlocke367):

but I'm confused.

OpenStudy (bloomlocke367):

if I had a graphing calculator this would be easy XD

OpenStudy (rational):

lets work 2-3 quick examples

OpenStudy (rational):

\[\begin{vmatrix} 1&2&3\\4&5&6\\7&8&9\end{vmatrix} =? \]

OpenStudy (rational):

there will be 3 terms in the determinant to get the first term ignore first row and first column, calculate the determinant of remaining 2x2 matrix : \[\begin{vmatrix} \color{gray}{1}& \color{gray}{2}& \color{gray}{3}\\ \color{gray}{4}&5&6\\ \color{gray}{7}&8&9\end{vmatrix} \]

OpenStudy (rational):

can you find the determinant of that bottom right 2x2 matrix ?

OpenStudy (bloomlocke367):

yes, -3?

OpenStudy (rational):

Yes! we call "-3" the cofactor of "1"

OpenStudy (bloomlocke367):

why?

OpenStudy (rational):

the first term in the determinant would be \(1(-3) = -3\)

OpenStudy (rational):

save that and lets find the second term, is it okay if i answer that question after we find the determinant ?

OpenStudy (bloomlocke367):

sure

OpenStudy (bloomlocke367):

how do I find the second term?

OpenStudy (rational):

to get the second term ignore first row and second column, calculate the determinant of remaining 2x2 matrix : \[\begin{vmatrix} \color{gray}{1}& \color{gray}{2}& \color{gray}{3}\\ 4&\color{gray}{5}&6\\ 7&\color{gray}{8}&9\end{vmatrix} \]

OpenStudy (bloomlocke367):

row, being "1 2 & 3" right?

OpenStudy (bloomlocke367):

just making sure I didn't get it confused..

OpenStudy (rational):

yes

OpenStudy (bloomlocke367):

rows are horizontal, and columns vertical, right?

OpenStudy (rational):

we're on same page

OpenStudy (bloomlocke367):

-6?

OpenStudy (rational):

Yes. so the second term would be \(-2(-6) = 12\)

OpenStudy (rational):

save that and lets find the third term

OpenStudy (bloomlocke367):

let me guess, first row, third column?

OpenStudy (bloomlocke367):

and why is it -2?

OpenStudy (rational):

to get the third term ignore first row and third column, calculate the determinant of remaining 2x2 matrix : \[\begin{vmatrix} \color{gray}{1}& \color{gray}{2}& \color{gray}{3}\\ 4&5&\color{gray}{6}\\ 7&8&\color{gray}{9}\end{vmatrix} \]

OpenStudy (bloomlocke367):

-3

OpenStudy (rational):

Yes so the third term would be \(3(-3) = -9\)

OpenStudy (rational):

Add these 3 terms to get the determinant

OpenStudy (rational):

determinant = \(-3 + 12 - 9 = ?\)

OpenStudy (bloomlocke367):

so why was it 1, -2, and 3? why was 2 negative?

OpenStudy (bloomlocke367):

and the answer is 0

OpenStudy (rational):

thats a good question! the sign is obtained using the formula : \(\large (-1)^{m+n}\) where m = row number n = column number

OpenStudy (rational):

you will be good with determinants if you just remember this : flip the sign of second term

OpenStudy (bloomlocke367):

okay...

OpenStudy (bloomlocke367):

so will you check this for me?

OpenStudy (rational):

sure ask

OpenStudy (bloomlocke367):

one sec, still working it out

OpenStudy (bloomlocke367):

how do you make a 3 x 3 matrix on here? XD

OpenStudy (rational):

``` \begin{vmatrix} a&b&c\\ d&e&f\\ g&h&i \end{vmatrix} ``` produces below \[\begin{vmatrix} a&b&c\\d&e&f\\g&h&i \end{vmatrix}\]

OpenStudy (bloomlocke367):

okay thanks

OpenStudy (bloomlocke367):

\(\begin{vmatrix}-4&5&6\\0&4&4\\-2&-5&4\end{vmatrix}\)

OpenStudy (bloomlocke367):

i got -20

OpenStudy (rational):

show steps for the three terms

OpenStudy (bloomlocke367):

okay. 1(-4) -2(-4) 3(-8) -4+8-24=-20

OpenStudy (bloomlocke367):

@rational

OpenStudy (rational):

NO

OpenStudy (bloomlocke367):

oh...

OpenStudy (rational):

\[\begin{vmatrix}-4&5&6\\0&4&4\\-2&-5&4\end{vmatrix} = -4(16+20) -5(0+8) +6(0+8) = ?\]

OpenStudy (bloomlocke367):

isn't it 16+-20? and 0+-8? and, again 0+-8?

OpenStudy (bloomlocke367):

and why is it 4, 5, and 6?

OpenStudy (rational):

determinant of 2x2 matrix is ad-bc right

OpenStudy (bloomlocke367):

yes.

OpenStudy (rational):

whats the determinant of | 4 4 | |-5 4 | ?

OpenStudy (bloomlocke367):

-4

OpenStudy (bloomlocke367):

wait, no

OpenStudy (bloomlocke367):

wow, I've been adding this entire time XD

OpenStudy (bloomlocke367):

thanks for bringing that to my attention. 16--20=16+20=36

OpenStudy (bloomlocke367):

so why is it 4, 5, and 6?

OpenStudy (rational):

look at first row

OpenStudy (bloomlocke367):

ohhh okay.

OpenStudy (bloomlocke367):

44?

OpenStudy (bloomlocke367):

what did I do wrong?...

OpenStudy (rational):

\[\begin{vmatrix}-4&5&6\\0&4&4\\-2&-5&4\end{vmatrix} = -4(16+20) -5(0+8) +6(0+8) = -4(36) - 40 + 48 = ?\]

OpenStudy (bloomlocke367):

I see what I did wrong xp I forgot to multiply -4 and 36 XD

OpenStudy (bloomlocke367):

well, I did, but when I did the last step, I totally forgot to put -144 in and put 36

OpenStudy (bloomlocke367):

now I got -136 :)

OpenStudy (bloomlocke367):

thank you

OpenStudy (rational):

nice :)